# To curve, or not to curve: Is curvature-assisted quintessence observationally viable?

### v1

@@ 5 Conclusions
In this paper we have studied the question of whether a nonzero (and specifically, negative) spatial curvature would make single-field quintessence models of cosmic acceleration with steep potentials observationally viable—this has been motivated by the results of Ref. Andriot et al. (2023).
We have first explained the importance of scalar fields in string theory, and after reviewing the theoretical reasons for why asymptotic corners of moduli space of string theoretic EFTs are interesting, we have argued that neither a de Sitter vacuum nor a quintessence scenario with a nearly flat potential can be realized in those corners. The reason is that all known asymptotic string theory examples seem to imply that there is a lower bound on the gradient of the potential that is too steep to provide accelerated expansion consistent with cosmological observations—this has led to a conjecture that this steepness is a universal feature of quantum gravity in the asymptotic regions of moduli space, and therefore, one may need to focus instead on the bulk of the space. However, the authors of the recent work Andriot et al. (2023) have argued that one could obtain observationally viable accelerated expansion even for steep potentials, i.e., from a fully controlled string theoretic EFT, if the Universe is negatively curved.
@@
The observational $1\sigma$ upper bound we have found on the quantity $\lambda$ is $\sim 0.6$ (consistent with the results of Refs. Agrawal et al. (2018); Akrami et al. (2019); Raveri et al. (2019)) for a flat universe, where the curvature density parameter $\Omega_{k}$ is fixed to zero, while we have found the $1\sigma$ upper bound of $\sim 0.7$ for the case in which $\Omega_{k}$ is varied as a free parameter, i.e., the curvature of the Universe is allowed to take nonzero values. Even though the curvature-assisted model allows $\lambda$ to take larger values compared to the zero-curvature model, the increase in the allowed values of $\lambda$ is too small to make the observational constraints consistent with the theoretical lower bound of $\sqrt{2}$ on $\lambda$. These results are consistent with our investigation of the phase space of the model, where we have performed a numerical grid analysis and shown that even though a new fixed point exists in the phase space when $\Omega_{k}>0$, none of the solutions of the set of differential equations describing the dynamical system simultaneously satisfy the two necessary conditions for a viable cosmology: 1) contain an extended period of matter domination before the onset of late-time cosmic acceleration, and 2) have an effective equation of state parameter that is $\sim-0.7$ at the present time.

### v2

@@ 5 Conclusions
In this paper we have studied the question of whether a nonzero (and specifically, negative) spatial curvature would make single-field quintessence models of cosmic acceleration with steep potentials observationally viable—this has been motivated by the results of Refs. Marconnet and Tsimpis (2023); Andriot et al. (2023).
We have first explained the importance of scalar fields in string theory, and after reviewing the theoretical reasons for why asymptotic corners of moduli space of string theoretic EFTs are interesting, we have argued that neither a de Sitter vacuum nor a quintessence scenario with a nearly flat potential can be realized in those corners. The reason is that all known asymptotic string theory examples seem to imply that there is a lower bound on the gradient of the potential that is too steep to provide accelerated expansion consistent with cosmological observations—this has led to a conjecture that this steepness is a universal feature of quantum gravity in the asymptotic regions of moduli space, and therefore, one may need to focus instead on the bulk of the space. However, the authors of the recent work Marconnet and Tsimpis (2023); Andriot et al. (2023) have argued that one could obtain observationally viable accelerated expansion even for steep potentials, i.e., from a fully controlled string theoretic EFT, if the Universe is negatively curved.
@@
The observational $1\sigma$ upper bound we have found on the quantity $\lambda$ is $\sim 0.6$ (consistent with the results of Refs. Agrawal et al. (2018); Akrami et al. (2019); Raveri et al. (2019)) for a flat universe, where the curvature density parameter $\Omega_{k}$ is fixed to zero, while we have found the $1\sigma$ upper bound of $\sim 0.7$ for the case in which $\Omega_{k,0}$ (the current value of the curvature density parameter) is varied as a free parameter, i.e., the curvature of the Universe is allowed to take nonzero values. Even though the curvature-assisted model allows $\lambda$ to take larger values compared to the zero-curvature model, the increase in the allowed values of $\lambda$ is too small to make the observational constraints consistent with the theoretical lower bound of $\sqrt{2}$ on $\lambda$. We have shown that our results are practically the same whether we fix the initial velocity of the scalar field $\dot{\phi}_{0}$ to zero or we set it free.
These results are consistent with our investigation of the phase space of the model (also allowing for an arbitrary value of the initial velocity of the scalar field), where we have performed a numerical grid analysis and shown that even though a new fixed point exists in the phase space when $\Omega_{k}>0$, none of the solutions of the set of differential equations describing the dynamical system simultaneously satisfy the two necessary conditions for a viable cosmology: 1) contain an extended period of matter domination before the onset of late-time cosmic acceleration, and 2) have an effective equation of state parameter that is $\sim-0.7$ at the present time.